OpenStudy (anonymous):

solve the exponential equation by expressing each side as a power of the same base and then equating exponents 6^x=36 the solution set is?

OpenStudy (anonymous):

\[6^x = 6^2\] x = 2

OpenStudy (anonymous):

on my math lab it says type in the solution set so what do i type in

OpenStudy (anonymous):

i guess it would be {2} but im not sure for that question